### v1

# Propagation and emission of gravitational waves in the weak-field limit within the Palatini formalism

Abstract
In the era of gravitational waves physics, when detections of wave fronts are increasing in number, sensibility, frequencies and distances, gravitational physics has entered a period of maximum activity and brilliance. This has open a new window where General Relativity can be challenged in both weak as strong-field regimes. In this paper, we focus on the analysis of gravitational waves propagation and emission in the weak-field regime for gravitational theories within the Palatini formalism. Our results suggest that while propagation in vacuum match General Relativity predictions, emission of gravitational waves from weak sources shows a clear deviation that might be constrained by future detections.
I Introduction
Over the last years, the field of gravitational waves (GWs) has experienced a great growth, mainly pushed forward by the many detections of GWs fronts by the interferometers of LIGO/VIRGO (for some of the first detections, see Refs. [1]). Such detections were the product of the coalescence of binary black holes systems as well as neutron stars, that allow also to detect the electromagnetic counterpart [2, 3]. Recently, the collaboration KAGRA has joined to network, which will increase the distance of the events and the number of coalescence binary neutron stars systems. In the next incoming years, other detectors such as LISA or later the Einstein Telescope will start operations where the sensibility and the range of frequencies will be increased allowing to detect sources at much higher redshifts and also GWs’ emission during the inspiral phase of some binary systems (see Refs. [4, 5]). Hence, one can say that one of the areas where gravitational physics will be focused for the next decades is GWs physics. From the theoretical point of view, this is a great opportunity to test the limits of General Relativity (GR). As has been widely analysed, GR has been an incredible successful theory despite also some limitations have shown up. From cosmology to compact objects physics, GR lacks of providing some convince explanations. One may cite dark energy, black holes singularities or the absence of ultraviolet completion of the theory, among others. Hence, along the next years GR will be tested as never has been before and compact objects physics will play an essential role, mainly through the analysis of GWs emissions [6]. Moreover, the incoming observational data at the cosmological level will also provide another independent source to test GR [7].
Over the last decades, modifications/extensions of GR have been widely studied, motivated by different phenomena, such analysis has led to a huge increase of knowledge on the way GR might be modified successfully (for a review see Refs. [8, 9, 10]). One of the ways of modifying GR that has been widely explored in the literature is to consider the connection as an independent field [11, 12, 13]. Despite GR is constructed as a geometrical theory that describes pseudo-Riemannian manifolds that are completely characterised by the spacetime metric, non of its principles states that the connection must be the Levi-Civita connection, i.e. that the metricity condition must hold. In addition, GR might be formulated in two different and physically equivalent ways, where the central magnitude is not the curvature but the torsion or the non-metricity [14]. For such theories, known as Teleparallel gravity and Coincident General Relativity respectively, the connection is not given by the Levi-Civita connection, but it is neither let as a free field. However, this has shown that GR might be modified in very different ways. In this paper, we focus on a particular case of the so-called Ricci Based Gravity (RBG) theories, whose starting point is a gravitational action that depends solely on contractions and products of the Ricci tensor, which is constructed by an a priori independent connection [12, 13]. By following the Palatini formalism, which consists on varying the action with respect to the connection as an independent field, the corresponding field equations are obtained for the connection. As far as the gravitational action turns out the Hilbert-Einstein action, the connection becomes the Levi-Civita connection of GR [15]. One of the main advantages of RBG theories is that the field equations for the spacetime metric remain second order while the equations for the connection leads to a direct solution that establishes a connection compatible to a metric tensor. Then, a particular mapping has been established that leads to the same equations as in GR, but now in terms of other metric tensor related to the spacetime metric a conformal transformation [16].The simplest version of RBG theories is the so-called Palatini $f(R)$ gravity where the action is given by a function of the contraction of the Ricci tensor scalar. The main advantage of Palatini $f(R)$ gravity is that the field equations for the spacetime metric remain second order while the equations for the connection leads to a direct solution that establishes a connection compatible to a metric tensor. These theories have been widely analysed in different frameworks, particularly a big effort has been followed for analysing the cosmological implications within these theories, as cosmological late-time acceleration [17, 18, 19, 20, 21] or inflationary models [22, 23, 24, 25, 26, 27]. Moreover, an intensive study has drawn a lot of attention in solutions of compact objects, since one can construct solutions for regular black holes/wormholes [28, 29, 30, 31, 32] or describe well stellar interiors [33, 34].
The present paper intends to focus on the analysis of GWs propagation and emission in the weak-field regime within Palatini $f(R)$ gravity. In general, scalar-tensor theories such as metric $f(R)$ gravities, lead to additional propagating scalar modes that yield to an additional longitudinal polarisation [35, 36, 37, 38, 39, 40, 41, 42]. However, some modifications of GR, such as the Palatini formalism itself [43, 44, 45] or some extensions of Teleparallel gravity [46] do not contain extra propagating modes. The analysis of GWs in modified gravities is essential to test the limits of GR and finding out the way GR might be extended. In this sense, some tests have been suggested in the literature, specifically for metric $f(R)$ gravities, where constraints on the free parameters of the models can be obtained [47, 48]. Moreover, also along the cosmological evolution some modifications of GR might affect some stochastic GWs backgrounds, specially during early cosmology [49, 50, 51]. In order to focus our analysis on the study of GWs in the weak-field limit for Palatini $f(R)$ gravity, we express this theory as a scalar-tensor theory (Brans-Dicke-like model). GWs have been widely explored within Brans-Dicke theories, both in the strong-field regime [52] as in the weak-field limit (see Refs. [54, 55, 56, 57, 58]). Nevertheless, contrary to usual Brans-Dicke-like theories, in Palatini $f(R)$ gravities, the scalar field is not dynamical, what has led the believe that GWs propagation and generation coincide with GR predictions. Here, we show that while propagating GWs in vacuum coincide with GR, emission of GWs has subtle differences with respect to GR, even in the weak field limit. We also point a way such deviations might be constrained.
The paper is organised as follows: modified Palatini $f(R)$ gravity is introduced in section II. Section III is devoted to the linearisation of the equations. Then, the propagation GWs equation in vacuum is obtained in section IV. Whereas the emission of GWs in the weak-field limit is analysed in section V. Finally, the last section VI gathers the conclusions of the paper.
II Modified Palatini $f(R)$ gravity
Along this paper we focus in the so-called Palatini $f(R)$ gravity, where the spacetime metric and the connection are considered as independent fields. The gravitational action is given by:
[…]
where $L_{m}$ is the Lagrangian that describes the matter fields and does not depend on the connection. The Ricci tensor is provided in terms of the independent connection as follows:
$\mathcal{R}_{\mu\nu}(\Gamma)=\partial_{\lambda}\Gamma^{\lambda}_{\mu\nu}-\partial_{\nu}\Gamma^{\lambda}_{\mu\lambda}+\Gamma^{\lambda}_{\sigma\lambda}\Gamma^{\sigma}_{\mu\nu}-\Gamma^{\lambda}_{\sigma\nu}\Gamma^{\sigma}_{\mu\lambda}\ .$ (2)
Whereas the curvature scalar for the gravitational action is $\mathcal{R}=g^{\mu\nu}\mathcal{R}_{\mu\nu}(\Gamma)$. Note that for the Hilbert-Einstein action, the Palatini formalism directly leads to GR, since the metricity condition is recovered, as shown below. However, for a non-linear function of $\mathcal{R}$ in (1), the metricity condition does not hold and the connection remains independent in principle. Then, by varying the action (1) with respect to the spacetime metric, the field equations are obtained [11]:
$f_{\mathcal{R}}\mathcal{R}_{\mu\nu}-\frac{1}{2}g_{\mu\nu}f=\kappa^{2}T_{\mu\nu}\ ,$ (3)
where $T_{\mu\nu}=-\frac{2}{\sqrt{-g}}\frac{\delta(\sqrt{-g}L_{m})}{\delta g^{\mu\nu}}$ is the energy-momentum tensor and $f_{\mathcal{R}}=\frac{df}{d\mathcal{R}}$. In addition, variations of the Ricci tensor with respect to the connection in the action lead to:
~$\tilde{\nabla}_{\lambda}\left(\sqrt{-g}f_{\mathcal{R}}g^{\mu\nu}\right)=0\ .$ (4)
Here $\tilde{\nabla}$ is the covariant derivative defined by the connection $\Gamma$. This equation gives the solution for the connection $\Gamma$ in terms of the Levi-Civita connection for the metric tensor:
$q_{\mu\nu}=\Omega^{2}g_{\mu\nu}\ ,\quad\Omega^{2}=f_{\mathcal{R}}\ ,$ (5)
Hence, one has ~$\tilde{\nabla}_{\lambda}\left(\sqrt{-q}q^{\mu\nu}\right)=0$, which reproduces the metricity condition for the covariant derivative $\tilde{\nabla}$ for the metric $q_{\mu\nu}$. Moreover, by taking the trace of the field equations (3), an algebraic relation between the scalar $\mathcal{R}$ and the energy-momentum tensor $T$ is obtained:
$f_{\mathcal{R}}\mathcal{R}-2f=\kappa^{2}T\ \quad\rightarrow\quad\mathcal{R}=\mathcal{R}(T)\ .$ (6)
By using the conformal relation given in (5), the field equations (3) may be rewritten in terms of the spacetime metric, instead of the independent connection. To do so, one has to make a conformal transformation of the Ricci tensor as follows:
$\mathcal{R}_{\mu\nu}(q)=R_{\mu\nu}(g)+\frac{4}{\Omega^{2}}\nabla_{\mu}\Omega\nabla_{\nu}\Omega-\frac{2}{\Omega}\nabla_{\mu}\nabla_{\nu}\Omega-g_{\mu\nu}\frac{g^{\rho\sigma}}{\Omega^{2}}\nabla_{\rho}\Omega\nabla_{\sigma}\Omega-g_{\mu\nu}\frac{\Box\Omega}{\Omega}\ .$ (7)
Now the covariant derivatives shown in the r.h.s. of (7) are compatible with the spacetime metric $g_{\mu\nu}$. Hence, the field equations (3) can be expressed as:
$R_{\mu\nu}(g)-\frac{1}{2}g_{\mu\nu}R(g)=\frac{\kappa^{2}}{f_{\mathcal{R}}}T_{\mu\nu}-g_{\mu\nu}\frac{\mathcal{R}f_{\mathcal{R}}-f}{2f_{\mathcal{R}}}-\frac{3}{2f_{\mathcal{R}}^{2}}\left[\nabla_{\mu}f_{\mathcal{R}}\nabla_{\nu}f_{\mathcal{R}}-\frac{1}{2}g_{\nu\mu}\nabla_{\lambda}f_{\mathcal{R}}\nabla^{\lambda}f_{\mathcal{R}}\right]+\frac{1}{f_{\mathcal{R}}}\left[\nabla_{\mu}\nabla_{\nu}f_{\mathcal{R}}-g_{\mu\nu}\Box f_{\mathcal{R}}\right]\ .$ (8)
In addition, it is straightforward to show that the field equations (8) are equivalent to the those of a scalar-tensor theory by defining the following scalar field $\phi$:
$\phi=f_{\mathcal{R}},\quad\quad\quad V(\phi)=\mathcal{R}\phi-f(\mathcal{R})\ .$ (9)
Then, the field equations yield:
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=\frac{\kappa^{2}}{\phi}T_{\mu\nu}-\frac{V(\phi)}{2\phi}g_{\mu\nu}-\frac{3}{2\phi^{2}}\left(\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}g_{\mu\nu}(\partial\phi)^{2}\right)+\frac{1}{\phi}\left(\nabla_{\mu}\nabla_{\nu}\phi-g_{\mu\nu}\Box\phi\right)\ ,$ (10)
Note also that the trace equation (6) can be written also in terms of the scalar field as:
$2V(\phi)-\phi V^{\prime}(\phi)=\kappa^{2}T\ ,$ (11)
[…]
III Linearisation of the field equations
Let us consider a small perturbation on the metric, on the scalar field and on the matter energy-momentum tensor around a particular background solution of the field equations:
$g_{\mu\nu}=\overline{g}_{\mu\nu}+h_{\mu\nu},\quad T_{\mu\nu}=\overline{T}_{\mu\nu}+\delta T_{\mu\nu},\quad\phi=\overline{\phi}+\delta\phi\ ,$ (12)
where the over-line is used to denote the background quantities, while $\delta T_{\mu\nu}$ and $\delta\phi$ are of the same order than the perturbation $h_{\mu\nu}$. The Einstein tensor at the l.h.s. of Eq. (10) can be expanded in the following standard way:
$G_{\mu\nu}$ $=R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R$
$=\overline{G}_{\mu\nu}+\frac{1}{2}\left(\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}h+\Box h\overline{g}_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\overline{g}_{\mu\nu}-\overline{R}h_{\mu\nu}\right),$ (13)
where $h=g^{\mu\nu}h_{\mu\nu}$, which at first order reads $h=\overline{g}^{\mu\nu}h_{\mu\nu}$. The covariant derivatives and D’Alembertians are also computed by using the background metric. Whereas the r.h.s. of field Eqs. (10) at first order in perturbations lead to:
$\kappa^{2}\frac{T_{\mu\nu}}{\overline{\phi}}+\kappa^{2}\frac{\delta T_{\mu\nu}}{\overline{\phi}}-\kappa^{2}\frac{T_{\mu\nu}}{\overline{\phi}}\frac{\delta\phi}{\overline{\phi}}-\frac{V(\overline{\phi})}{2\overline{\phi}}\overline{g}_{\mu\nu}+\frac{V(\overline{\phi})}{2\overline{\phi}}h_{\mu\nu}+\frac{V(\overline{\phi})}{2\overline{\phi}}\overline{g}_{\mu\nu}\frac{\delta\phi}{\overline{\phi}}$
$-\left(\frac{3}{2\overline{\phi}^{2}}-\frac{3}{\overline{\phi}^{2}}\frac{\delta\phi}{\overline{\phi}}\right)\left[\partial_{\mu}\overline{\phi}\partial_{\nu}\overline{\phi}+\partial_{\mu}\overline{\phi}\partial_{\nu}\delta\phi+\partial_{\mu}\delta\phi\partial_{\nu}\overline{\phi}-\frac{1}{2}\overline{g}_{\nu\mu}(\partial\overline{\phi})^{2}-\frac{1}{2}h_{\mu\nu}(\partial\overline{\phi})^{2}-\overline{g}_{\mu\nu}\partial^{\alpha}\partial_{\alpha}\delta\phi\right]$
$+\frac{1}{\overline{\phi}}\Bigg{[}\nabla_{\mu}\nabla_{\nu}\overline{\phi}+\nabla_{\mu}\nabla_{\nu}\delta\phi-\frac{\overline{g}^{\lambda\alpha}}{2}(\nabla_{\nu}h_{\alpha\mu}+\nabla_{\mu}h_{\alpha\nu}-\nabla_{\alpha}h_{\mu\nu})\partial_{\lambda}\overline{\phi}-\overline{g}_{\mu\nu}\Box\overline{\phi}-\overline{g}_{\mu\nu}\Box\delta\phi$
$+\overline{g}_{\mu\nu}\frac{\overline{g}^{\lambda\alpha}}{2}(\nabla_{\beta}h_{\lambda\alpha}+\nabla_{\alpha}h_{\lambda\beta}-\nabla_{\lambda}h_{\alpha\beta})-h_{\mu\nu}\Box\overline{\phi}-\nabla_{\mu}\nabla_{\nu}\overline{\phi}\frac{\delta\phi}{\overline{\phi}}-\overline{g}_{\mu\nu}\Box\overline{\phi}\frac{\delta\phi}{\overline{\phi}}\Bigg{]}.$ (14)
Hence, by taking both sides, finally the field equations at first order are obtained:
$\frac{1}{2}\left(\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}h+\Box h\overline{g}_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\overline{g}_{\mu\nu}-\overline{R}h_{\mu\nu}\right)=$
$=\kappa^{2}\frac{\delta T_{\mu\nu}}{\overline{\phi}}-\kappa^{2}\frac{T_{\mu\nu}}{\overline{\phi}}\frac{\delta\phi}{\overline{\phi}}+\frac{V(\overline{\phi})}{2\overline{\phi}}h_{\mu\nu}+\frac{V(\overline{\phi})}{2\overline{\phi}}\overline{g}_{\mu\nu}\frac{\delta\phi}{\overline{\phi}}$
$-\frac{3}{2\overline{\phi}^{2}}\left[\partial_{\mu}\overline{\phi}\partial_{\nu}\delta\phi+\partial_{\mu}\delta\phi\partial_{\nu}\overline{\phi}-\frac{1}{2}h_{\mu\nu}(\partial\overline{\phi})^{2}-\overline{g}_{\mu\nu}\partial^{\alpha}\partial_{\alpha}\delta\phi\right]+\frac{3}{\overline{\phi}^{2}}\frac{\delta\phi}{\overline{\phi}}\left[\partial_{\mu}\overline{\phi}\partial_{\nu}\overline{\phi}-\frac{1}{2}\overline{g}_{\nu\mu}(\partial\overline{\phi})^{2}\right]$
$+\frac{1}{\overline{\phi}}\Bigg{[}\nabla_{\mu}\nabla_{\nu}\delta\phi-\frac{\overline{g}^{\lambda\alpha}}{2}(\nabla_{\nu}h_{\alpha\mu}+\nabla_{\mu}h_{\alpha\nu}-\nabla_{\alpha}h_{\mu\nu})\partial_{\lambda}\overline{\phi}-\overline{g}_{\mu\nu}\Box\delta\phi$
$+\overline{g}_{\mu\nu}\frac{\overline{g}^{\lambda\alpha}}{2}(\nabla_{\beta}h_{\lambda\alpha}+\nabla_{\alpha}h_{\lambda\beta}-\nabla_{\lambda}h_{\alpha\beta})-h_{\mu\nu}\Box\overline{\phi}-\nabla_{\mu}\nabla_{\nu}\overline{\phi}\frac{\delta\phi}{\overline{\phi}}-\overline{g}_{\mu\nu}\Box\overline{\phi}\frac{\delta\phi}{\overline{\phi}}\Bigg{]}.$ (15)
In addition, the scalar field equation at first order (11) turns out:
$\left(V^{\prime}(\overline{\phi})-\overline{\phi}V^{\prime\prime}(\overline{\phi})\right)\delta\phi=\kappa^{2}\delta T.$ (16)
From here, we can now analyse the propagation of gravitational waves in vacuum and also the generation of them in the weak field limit.
IV Propagation of gravitational waves in vacuum
While the linearisation of the field equations can be used to analyse more complex scenarios, as black hole perturbations and the corresponding stability, here we aim to focus on the simple aspect of the propagation in vacuum of the GW. Hence, for the weak field limit and by considering the background spacetime metric as the Minkowski metric, one has:
$\overline{R}=0\ ,\quad T_{\mu\nu}=0\ ,\quad\overline{\phi}=\phi_{0}\ .$ (17)
Then, by definition, $V^{\prime}(\phi_{0})=0$, and in order to satisfy the background solution $V(\phi_{0})=0$ must hold. Hence, the equation for the perturbations (15) reads:
$\frac{1}{2}\left(\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}h+\Box h\eta_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\eta_{\mu\nu}\right)=\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}-\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ .$ (18)
This equation is free from gauge choices and covariant derivatives are actually partial derivatives, as correspond for a flat background metric. In order to simplify the equations (18), we can choose the Lorentz Gauge $\partial^{\alpha}h_{\alpha\beta}=0$ to obtain a wave equation. Note also that the r.h.s. of the equations (18) automatically satisfies the Lorentz gauge. By defining the new tensor $\theta_{\mu\nu}=h_{\mu\nu}-\frac{1}{2}h\eta_{\mu\nu}$ the gravitational wave equation yields:
$\Box\theta_{\mu\nu}=-\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}+\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ .$ (19)
In comparison to GR, now the GWs equations in vacuum are not effectively in vacuum but are sourced by the perturbation of the scalar field. To show that actually such a perturbation does not play any role (in vacuum), let’s see that in vacuum Eq. (16) implies:
$-\phi_{0}V^{\prime\prime}(\phi_{0})\delta\phi=0\ ,$ (20)
However, this equation may imply three different scenarios depending actually on the theory:
  • For $\phi_{0}\neq 0$ and $V^{\prime\prime}(\phi_{0})\neq 0$, the perturbation for the scalar field becomes null $\delta\phi=0$ and consequently, the GW equation in vacuum turns out the same as in GR.
  • For $\phi_{0}=0$. This would imply that that $\phi_{0}=f_{\mathcal{R}_{0}}=0$ and consequently the field equations (8) or (10) are ill-defined for a Minkowski background, which does not correspond to a viable theory.
  • $V^{\prime\prime}(\phi_{0})=0$. This is probably the most interesting case, as it might be in principle possible as far as the scalar field potential and consequently the gravitational action $f_{\mathcal{R}}$ satisfies such condition. However, for the vacuum case $T=0$, the non-perturbative equation (11) implies that the scalar field becomes constant, so the perturbation $\delta\phi$ turns out a constant as well, which implies that the GW equation (19) reduces again to the standard case of GR.
Hence, the only viable possibilities makes the GW equation (19) to become:
$\Box\theta_{\mu\nu}=0\ .$ (21)
This is the GW equation for GR in vacuum, such that GWs in the Palatini formalism behave the same way in vacuum as in GR, with the same speed of propagation $c$ and the same two types of polarisations. Note that this is not the case for most of scalar-tensor theories, including metric $f(R)$ gravity, where some new longitudinal polarisations arise and/or the waves do not propagate at the speed of light. However, in the next section, we show that even in the weak field limit, when sources provided by an energy-momentum tensor are present, solutions become different as in GR.
V Emission of gravitational waves in the weak-field limit
Our aim here is to show the implications for the generation of GWs in Palatini $f(R)$ theories by analysing the field equations in the weak limit. To do so, we follow a standard procedure by assuming the sources to be weak enough in order to consider the background spacetime metric as approximately flat, i.e. as the Minkowski metric while the energy-momentum tensor $T_{\mu\nu}$ is already a first order perturbation (for details on this procedure see for instance [59]). By assuming the Lorentz gauge, the equations for the GWs (15) reduce to:
$\Box\theta_{\mu\nu}=-\frac{2\kappa^{2}}{\phi_{0}}T_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}+\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ ,$ (22)
where remind that we have defined $\theta_{\mu\nu}=h_{\mu\nu}-\frac{1}{2}h\eta_{\mu\nu}$. While the equation that relates the scalar field perturbation with the energy-momentum tensor (16) becomes:
$-\phi_{0}V^{\prime\prime}(\phi_{0})\delta\phi=\kappa^{2}T\ .$ (23)
Note that now the scalar field perturbation depends directly on the source, specifically on the trace of the energy-momentum tensor, such that for traceless matter, as radiation, the perturbation turns out null and the usual GWs equation of GR are recovered again. However, besides such a case, the scalar field perturbation will not be in general null neither constant. In fact, by combining both equations (22) and (23), the situation becomes clearer:
$\Box\theta_{\mu\nu}=-\frac{2\kappa^{2}}{\phi_{0}}T_{\mu\nu}+\frac{1}{\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\left(\nabla_{\mu}\nabla_{\nu}T-\eta_{\mu\nu}\Box T\right)\ .$ (24)
Hence, the generation of the GWs does not depend only on the presence of perturbed sources of matter but also on the way the energy-momentum tensor varies, as denoted by the presence of second partial derivatives in the equation (24). However, one would expect that for matter sources varying very slowly, the last two terms of the equation (24) might be neglected, recovering the usual case of GR. Nevertheless, in order to keep the analysis as general as possible and study the possible deviations from GR, we are considering that the last two terms might be the same order as the energy-momentum tensor perturbation itself. To solve the equation (24), we define the effective energy momentum tensor as:
$T_{\mu\nu}^{\text{eff}}=\frac{1}{\phi_{0}}T_{\mu\nu}-\frac{1}{2\kappa^{2}\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\left(\nabla_{\mu}\nabla_{\nu}T-\eta_{\mu\nu}\Box T\right)\ ,$ (25)
Then, the GW equation (22) can be rewritten in the standard GR case as:
$\Box\theta_{\mu\nu}=-2\kappa^{2}T_{\mu\nu}^{\text{eff}}\ .$ (26)
This equation can be solved by following the Green function method, leading to the solution:
$\theta_{\mu\nu}=-2\kappa^{2}\int d^{4}x^{\prime}G(x-x^{\prime})T_{\mu\nu}^{\text{eff}}(x^{\prime})\ ,$ (27)
where the Green function is provided as far as satisfies the equation:
$\Box G(x-x^{\prime})=\delta^{4}(x-x^{\prime})\ .$ (28)
The solution of this equation is given by the retarded Green function:
$G(x-x^{\prime})=-\frac{1}{4\pi|\vec{x}-\vec{x}^{\prime}|}\delta(x_{ret}^{0}-x^{0\prime})\ .$ (29)
Here $\vec{x}$ are spatial vectors, $x^{0}=t$ and $x_{ret}^{0}=t_{ret}=t-|\vec{x}-\vec{x}^{\prime}|$ is the retarded time and we are assuming $c=1$. By considering an enough large distance $|\vec{x}|$ from the source, we can approximate $|\vec{x}-\vec{x}^{\prime}|=r-\vec{x}^{\prime}\cdot\hat{n}$, where $r=|\vec{x}|$ and $\hat{n}$ is the unit vector pointing to the direction of the GW propagation. For simplicity, we work in the transverse-traceless gauge, which by following the standard notation, it is given by:
$\theta_{ij}^{TT}=\Lambda_{ijkl}\theta^{kl}\ ,$ (30)
where $\theta_{kl}$ is the spatial part of $\theta_{\mu\nu}$, whose indexes are upper and lower by the flat metric $\eta_{ij}$, whereas the tensor $\Lambda_{ijkl}$ is:
$\Lambda_{ijkl}(\hat{n})=P_{ik}P_{jl}-\frac{1}{2}P_{ij}P_{kl}\ ,\quad P_{ij}=\delta_{ij}-n_{i}n_{j}\ .$ (31)
Then, by inserting the Green function (29) in the expression (27) and integrating over $x^{0\prime}$, it yields:
$\theta_{ij}^{TT}=\frac{4G}{r}\Lambda_{ijkl}\int d^{3}x^{\prime}T^{kl}_{\text{eff}}(t-r+\vec{x}^{\prime}\cdot\hat{n},x^{\prime})\ .$ (32)
The aim now is to perform a multipole expansion and show the dependence of the quadrupole moment for this class of theories. To do so, let’s expand the energy-momentum tensor. For enough large distance, we can make the following expansion:
$T_{kl}^{\text{eff}}(t-\frac{r}{c}+\frac{\vec{x}^{\prime}\cdot\hat{n}}{c},x^{\prime})=T_{kl}^{\text{eff}}(t-r,x^{\prime})+\frac{\partial T_{kl}^{\text{eff}}}{dt}\frac{\vec{x}^{\prime}\cdot\hat{n}}{c}+\frac{\partial^{2}T_{kl}^{\text{eff}}}{dt^{2}}\frac{(\vec{x}^{\prime}\cdot\hat{n})(\vec{x}^{\prime}\cdot\hat{n})}{2c^{2}}+\mathcal{O}\left(\frac{1}{c^{3}}\right)\ ,$ (33)
where we have recovered the speed of light to show that the expansion is inversely proportional to $c$. Hence, the expression (32) becomes:
$\theta_{ij}^{TT}=\frac{4G}{r}\Lambda_{ijkl}\left(S^{kl}+\frac{1}{c}n_{m}\dot{S}^{klm}+\frac{1}{2c^{2}}n_{m}n_{p}\ddot{S}^{klmp}+...\right)\ .$ (34)
Here we have defined the tensors $S^{kl..}$ as follows:
$S^{kl}=\int d^{3}x^{\prime}T_{\text{eff}}^{kl}\ ,\quad S^{klm}=\int d^{3}x^{\prime}T_{\text{eff}}^{kl}x^{m\prime}\ .$ (35)
Let us just focus on the leading term in (34). To write such a term in a more familiar way, we use the fact that:
$\partial_{\mu}T^{\mu\nu}_{\text{eff}}=0\ .$ (36)
Note also that this implies automatically the conservation of the energy-momentum tensor $\partial_{\mu}T^{\mu\nu}=0$, since the two remaining terms in (25) cancel each other when taking the divergence. Then, the leading term in (35) yield:
$S^{kl}$ $=$ $\int d^{3}x^{\prime}T_{\text{eff}}^{km}\delta_{m}^{l}=\int d^{3}x^{\prime}\frac{\partial}{\partial x^{m\prime}}\left(T_{\text{eff}}^{km}x^{l\prime}\right)-\int d^{3}x^{\prime}\frac{\partial T_{\text{eff}}^{km}}{\partial x^{m\prime}}x^{l\prime}=$ (37)
$=$ $-\int d^{3}x^{\prime}\frac{\partial T_{\text{eff}}^{km}}{\partial x^{m\prime}}x^{l\prime}=\int d^{3}x^{\prime}\dot{T}_{\text{eff}}^{k0}x^{l\prime}=\int d^{3}x^{\prime}\ddot{T}_{\text{eff}}^{00}x^{k\prime}x^{l\prime}\ .$ (38)
Here we have used the continuity equation (36) and the volume of integration is taken over the source, such that boundary terms are removed. Moreover, in the last step, the same procedure as previous steps is followed. Then, by taking now the expression for the effective energy-momentum tensor (25), the expression for the quadrupole moment tensor is finally obtained:
$S^{kl}=\frac{1}{\phi_{0}}\frac{\partial^{2}}{\partial t^{2}}\int d^{3}x^{\prime}\left(T^{00}-\frac{1}{\phi_{0}V_{0}^{\prime\prime}}\partial_{j}\partial^{j}T\right)x^{k\prime}x^{l\prime}=\ddot{Q}^{kl}\ ,$ (39)
where
$Q^{kl}=\frac{1}{\phi_{0}}\int d^{3}x^{\prime}\left(T^{00}-\frac{1}{\phi_{0}V_{0}^{\prime\prime}}\partial_{j}\partial^{j}T\right)x^{k\prime}x^{l\prime}$ (40)
is the quadrupole moment for Palatini $f(R)$ gravity. Then, the expression for the gravitational wave (34) at the quadrupole order yields:
$\left[\theta_{ij}^{TT}\right]_{\text{quad}}=\frac{4G}{rc^{4}}\Lambda_{ijkl}\ddot{Q}^{kl}\ .$ (41)
The first term in the r.h.s. of (40) corresponds to the usual one of GR but with an effective gravitational constant given by $G_{eff}=\frac{G}{\phi_{0}}$. Nevertheless, the second term turns out a new contribution that just shown up in this class of gravitational theories. Such contribution is removed for a homogeneous and isotropic energy-momentum tensor, but might become important in other cases. However, even whether the tensor $T^{\mu\nu}$ represents a inhomogeneous and/or anisotropic source, one would expect that variations along different spatial directions might be small enough to become negligible in the weak-field limit. Nevertheless, such a term depends on the vacuum expectation value of the scalar field potential and specifically on its second derivative, that can be defined as the scalar field effective mass:
$m_{eff}^{2}=V^{\prime\prime}(\phi_{0})\ .$ (42)
Hence, one can conclude that for massive scalar fields, the second contribution in (40) can be neglected while light fields might provide an important contribution to the generation of GWs, even whether the source is nearly homogeneous and isotropic but not completely. In addition, at the quadrupole order the energy radiated with the gravitational wave is:
$\left[\frac{dE}{dt}\right]_{\text{quad}}\propto\langle\ddot{Q}^{kl}\ddot{Q}^{kl}\rangle\ ,$ (43)
[…]
VI Conclusions
Along this paper, the analysis of GWs propagation and emission is carried out for $f(R)$ theories within the Palatini formalism. To do so, firstly the field equations are expressed in a scalar-tensor-like form, which lead to the presence of a non-dynamical scalar field but sourced by the presence of matter, according to the scalar field equation itself. Then, the full set of field equations are linearised at first order in perturbations of the spacetime metric and the scalar field over an arbitrary background. To analyse the weak-field regime, we focus then on studying first order tensor perturbations of the metric in vacuum and also with the presence of weak enough sources. For the former case, the usual wave equation of GR is recovered, such that GWs propagation within the Palatini formalism obey the same rules as in GR with respect to polariations and speed of propagation.
Nevertheless, when assuming the presence of sources and analysing the emission of GWs through retarded Green functions, one finds that the multipoles expansion not only depend on time variations of the energy density, as provided by the 00-component of the energy-momentum tensor but also in spatial derivatives of its trace. Hence, for homogeneous and isotropic sources, the same results as in GR are recovered and the lowest order in the multipole expansion (the quadrupole moment) leads to the dependence as in GR. However, the presence of inhomogeneous sources introduces inevitable corrections to every order at the expansion and particularly, it modifies the quadrupole moment, leading to a stronger or weaker GWS emission, depending on the source. We suggest that for an enough massive scalar field, such corrections are neglected, while light scalar fields are ruled out, since will introduce strong deviations, making the energy radiated by the source too weak or too strong to satisfy observational constraints.
Next steps suggest that GWs must be analysed as well in strong-field regimes within Palatini-like theories, since the analysis of this paper suggest that while in vacuum, usual perturbation equations of GR are recovered, this changes completely with the presence of sources. In addition, the extension of these analysis to arbitrary RBG theories will provide a grateful insight on the understanding of theories within the Palatini formalism.
Acknowledgments
This work is supported by the Spanish National Grant PID2020-117301GA-I00 (ADC and DSG) funded by MCIN/AEI/10.13039/501100011033 (“ERDF A way of making Europe” and “PGC Generación de Conocimiento”). ADC is also funded by a pre-doctoral contract from the Predoctoral Contracts UVa 2022 co-financed by Banco Santander.

### v2

@@
 In this paper, we focus on the analysis of gravitational waves propagation and emission in the weak-field regime for gravitational theories within the Palatini formalism.
-Our results suggest that while propagation in vacuum match General Relativity predictions, emission of gravitational waves from weak sources shows a clear deviation that might be constrained by future detections.
+Our results show that gravitational waves propagation in vacuum matches General Relativity predictions as well as the multipolar expansion when considering weak sources, independently of the form of the gravitational action.
 I Introduction
@@
 The main advantage of Palatini $f(R)$ gravity is that the field equations for the spacetime metric remain second order while the equations for the connection leads to a direct solution that establishes a connection compatible to a metric tensor.
-These theories have been widely analysed in different frameworks, particularly a big effort has been followed for analysing the cosmological implications within these theories, as cosmological late-time acceleration [17, 18, 19, 20, 21] or inflationary models [22, 23, 24, 25, 26, 27].
+These theories have been widely analysed in different frameworks, particularly a big effort has been followed for analysing the cosmological implications within these theories, as cosmological late-time acceleration [17, 18, 19, 20, 21] or inflationary models [22, 23, 24, 25, 26, 27, 28, 29].
 Moreover, an intensive study has drawn a lot of attention in solutions of compact objects, since one can construct solutions for regular black holes/wormholes [30, 31, 32, 33, 34] or describe well stellar interiors [35, 36].
@@
 However, some modifications of GR, such as the Palatini formalism itself [45, 46, 47] or some extensions of Teleparallel gravity [48] do not contain extra propagating modes.
-The analysis of GWs in modified gravities is essential to test the limits of GR and finding out the way GR might be extended.
-In this sense, some tests have been suggested in the literature, specifically for metric $f(R)$ gravities, where constraints on the free parameters of the models can be obtained [47, 48].
-Moreover, also along the cosmological evolution some modifications of GR might affect some stochastic GWs backgrounds, specially during early cosmology [49, 50, 51].
+So far, GWs properties in certain theories of modified gravities have been found by using the post-Newtonian formalism [49, 50, 51, 52] and the WKB approach [53].
+The analysis of GWs in modified gravities is essential to test the limits of GR and to find out the way GR might be extended.
+In this sense, some tests have been suggested in the literature, specially for metric $f(R)$ gravities where constraints on the free parameters of the models can be obtained [54, 55].
+Moreover, some modifications of GR might affect stochastic GWs backgrounds, specially during early cosmology [56, 57, 58].
 In order to focus our analysis on the study of GWs in the weak-field limit for Palatini $f(R)$ gravity, we express this theory as a scalar-tensor theory (Brans-Dicke-like model).
@@
 [61, 62, 63, 64, 65]).
-Nevertheless, contrary to usual Brans-Dicke-like theories, in Palatini $f(R)$ gravities, the scalar field is not dynamical, what has led the believe that GWs propagation and generation coincide with GR predictions.
-Here, we show that while propagating GWs in vacuum coincide with GR, emission of GWs has subtle differences with respect to GR, even in the weak field limit.
-We also point a way such deviations might be constrained.
+Nevertheless, contrary to usual Brans-Dicke-like theories, the scalar field is not dynamical in Palatini $f(R)$ gravities.
+Then, by following such procedure we show that propagating GWs in vacuum coincide with GR predictions, as shown previously in [53] by following a different approach.
+Moreover, by considering weak sources, the corresponding multipole expansion is performed, where the quadrupole and octupole moments are computed, showing that despite the GWs equation is sourced by an effective energy-momentum tensor that includes derivatives of the trace of the energy-momentum tensor, each order of the multipole expansion reduce to the ones predicted by GR.
 The paper is organised as follows: modified Palatini $f(R)$ gravity is introduced in section II.
@@
 Whereas the curvature scalar for the gravitational action is $\mathcal{R}=g^{\mu\nu}\mathcal{R}_{\mu\nu}(\Gamma)$.
+The independent connection is not necessarily torsionless a priori, however, only its symmetric part plays a role on the equations of motion (see.
+[11]).
 Note that for the Hilbert-Einstein action, the Palatini formalism directly leads to GR, since the metricity condition is recovered, as shown below.
@@
 To do so, one has to make a conformal transformation of the Ricci tensor as follows:
-$\mathcal{R}_{\mu\nu}(q)=R_{\mu\nu}(g)+\frac{4}{\Omega^{2}}\nabla_{\mu}\Omega\nabla_{\nu}\Omega-\frac{2}{\Omega}\nabla_{\mu}\nabla_{\nu}\Omega-g_{\mu\nu}\frac{g^{\rho\sigma}}{\Omega^{2}}\nabla_{\rho}\Omega\nabla_{\sigma}\Omega-g_{\mu\nu}\frac{\Box\Omega}{\Omega}\ .$ (7)
+$\mathcal{R}_{\mu\nu}(q)=R_{\mu\nu}(g)+\frac{4}{\Omega^{2}}\nabla_{\mu}\Omega\nabla_{\nu}\Omega-\frac{2}{\Omega}\nabla_{\mu}\nabla_{\nu}\Omega$
+$-g_{\mu\nu}\frac{g^{\rho\sigma}}{\Omega^{2}}\nabla_{\rho}\Omega\nabla_{\sigma}\Omega-g_{\mu\nu}\frac{\Box\Omega}{\Omega}\ .$ (7)
 Now the covariant derivatives shown in the r.h.s. of (7) are compatible with the spacetime metric $g_{\mu\nu}$. Hence, the field equations (3) can be expressed as:
-$R_{\mu\nu}(g)-\frac{1}{2}g_{\mu\nu}R(g)=\frac{\kappa^{2}}{f_{\mathcal{R}}}T_{\mu\nu}-g_{\mu\nu}\frac{\mathcal{R}f_{\mathcal{R}}-f}{2f_{\mathcal{R}}}-\frac{3}{2f_{\mathcal{R}}^{2}}\left[\nabla_{\mu}f_{\mathcal{R}}\nabla_{\nu}f_{\mathcal{R}}-\frac{1}{2}g_{\nu\mu}\nabla_{\lambda}f_{\mathcal{R}}\nabla^{\lambda}f_{\mathcal{R}}\right]+\frac{1}{f_{\mathcal{R}}}\left[\nabla_{\mu}\nabla_{\nu}f_{\mathcal{R}}-g_{\mu\nu}\Box f_{\mathcal{R}}\right]\ .$ (8)
+$R_{\mu\nu}(g)-\frac{1}{2}g_{\mu\nu}R(g)=\frac{\kappa^{2}}{f_{\mathcal{R}}}T_{\mu\nu}-g_{\mu\nu}\frac{\mathcal{R}f_{\mathcal{R}}-f}{2f_{\mathcal{R}}}$
+$-\frac{3}{2f_{\mathcal{R}}^{2}}\left[\nabla_{\mu}f_{\mathcal{R}}\nabla_{\nu}f_{\mathcal{R}}-\frac{1}{2}g_{\nu\mu}\nabla_{\lambda}f_{\mathcal{R}}\nabla^{\lambda}f_{\mathcal{R}}\right]$
+$+\frac{1}{f_{\mathcal{R}}}\left[\nabla_{\mu}\nabla_{\nu}f_{\mathcal{R}}-g_{\mu\nu}\Box f_{\mathcal{R}}\right]\ .$ (8)
 In addition, it is straightforward to show that the field equations (8) are equivalent to the those of a scalar-tensor theory by defining the following scalar field $\phi$:
@@
 Then, the field equations yield:
-$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=\frac{\kappa^{2}}{\phi}T_{\mu\nu}-\frac{V(\phi)}{2\phi}g_{\mu\nu}-\frac{3}{2\phi^{2}}\left(\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}g_{\mu\nu}(\partial\phi)^{2}\right)+\frac{1}{\phi}\left(\nabla_{\mu}\nabla_{\nu}\phi-g_{\mu\nu}\Box\phi\right)\ ,$ (10)
+$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=\frac{\kappa^{2}}{\phi}T_{\mu\nu}-\frac{V(\phi)}{2\phi}g_{\mu\nu}$
+$-\frac{3}{2\phi^{2}}\left(\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}g_{\mu\nu}(\partial\phi)^{2}\right)$
+$+\frac{1}{\phi}\left(\nabla_{\mu}\nabla_{\nu}\phi-g_{\mu\nu}\Box\phi\right)\ ,$ (10)
 Note also that the trace equation (6) can be written also in terms of the scalar field as:
@@
 $G_{\mu\nu}$ $=R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R$
-$=\overline{G}_{\mu\nu}+\frac{1}{2}\left(\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}h+\Box h\overline{g}_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\overline{g}_{\mu\nu}-\overline{R}h_{\mu\nu}\right),$ (13)
+$=\overline{G}_{\mu\nu}+\frac{1}{2}\big{(}\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}$
+$-\nabla_{\nu}\nabla_{\mu}h+\Box h\overline{g}_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\overline{g}_{\mu\nu}-\overline{R}h_{\mu\nu}$
+$+\overline{g}_{\mu\nu}h^{\alpha\beta}\overline{R}_{\alpha\beta}\big{)},$ (13)
 where $h=g^{\mu\nu}h_{\mu\nu}$, which at first order reads $h=\overline{g}^{\mu\nu}h_{\mu\nu}$.
@@
 Hence, by taking both sides, finally the field equations at first order are obtained:
-$\frac{1}{2}\left(\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}h+\Box h\overline{g}_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\overline{g}_{\mu\nu}-\overline{R}h_{\mu\nu}\right)=$
+$\frac{1}{2}\left(\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}-\nabla_{\nu}\nabla_{\mu}h+\Box h\overline{g}_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\overline{g}_{\mu\nu}-\overline{R}h_{\mu\nu}+\overline{g}_{\mu\nu}h^{\alpha\beta}\overline{R}_{\alpha\beta}\right)=$
 $=\kappa^{2}\frac{\delta T_{\mu\nu}}{\overline{\phi}}-\kappa^{2}\frac{T_{\mu\nu}}{\overline{\phi}}\frac{\delta\phi}{\overline{\phi}}+\frac{V(\overline{\phi})}{2\overline{\phi}}h_{\mu\nu}+\frac{V(\overline{\phi})}{2\overline{\phi}}\overline{g}_{\mu\nu}\frac{\delta\phi}{\overline{\phi}}$
-$-\frac{3}{2\overline{\phi}^{2}}\left[\partial_{\mu}\overline{\phi}\partial_{\nu}\delta\phi+\partial_{\mu}\delta\phi\partial_{\nu}\overline{\phi}-\frac{1}{2}h_{\mu\nu}(\partial\overline{\phi})^{2}-\overline{g}_{\mu\nu}\partial^{\alpha}\partial_{\alpha}\delta\phi\right]+\frac{3}{\overline{\phi}^{2}}\frac{\delta\phi}{\overline{\phi}}\left[\partial_{\mu}\overline{\phi}\partial_{\nu}\overline{\phi}-\frac{1}{2}\overline{g}_{\nu\mu}(\partial\overline{\phi})^{2}\right]$
+$-\frac{3}{2\overline{\phi}^{2}}\left[\partial_{\mu}\overline{\phi}\partial_{\nu}\delta\phi+\partial_{\mu}\delta\phi\partial_{\nu}\overline{\phi}-\frac{1}{2}h_{\mu\nu}(\partial\overline{\phi})^{2}-\overline{g}_{\mu\nu}\partial^{\alpha}\overline{\phi}\partial_{\alpha}\delta\phi\right]+\frac{3}{\overline{\phi}^{2}}\frac{\delta\phi}{\overline{\phi}}\left[\partial_{\mu}\overline{\phi}\partial_{\nu}\overline{\phi}-\frac{1}{2}\overline{g}_{\nu\mu}(\partial\overline{\phi})^{2}\right]$
 $+\frac{1}{\overline{\phi}}\Bigg{[}\nabla_{\mu}\nabla_{\nu}\delta\phi-\frac{\overline{g}^{\lambda\alpha}}{2}(\nabla_{\nu}h_{\alpha\mu}+\nabla_{\mu}h_{\alpha\nu}-\nabla_{\alpha}h_{\mu\nu})\partial_{\lambda}\overline{\phi}-\overline{g}_{\mu\nu}\Box\delta\phi$
-$+\overline{g}_{\mu\nu}\frac{\overline{g}^{\lambda\alpha}}{2}(\nabla_{\beta}h_{\lambda\alpha}+\nabla_{\alpha}h_{\lambda\beta}-\nabla_{\lambda}h_{\alpha\beta})-h_{\mu\nu}\Box\overline{\phi}-\nabla_{\mu}\nabla_{\nu}\overline{\phi}\frac{\delta\phi}{\overline{\phi}}-\overline{g}_{\mu\nu}\Box\overline{\phi}\frac{\delta\phi}{\overline{\phi}}\Bigg{]}.$ (15)
+$+\overline{g}_{\mu\nu}\frac{\overline{g}^{\lambda\alpha}}{2}(\nabla_{\beta}h_{\lambda\alpha}+\nabla_{\alpha}h_{\lambda\beta}-\nabla_{\lambda}h_{\alpha\beta})\partial_{\beta}\overline{\phi}-h_{\mu\nu}\Box\overline{\phi}-\nabla_{\mu}\nabla_{\nu}\overline{\phi}\frac{\delta\phi}{\overline{\phi}}-\overline{g}_{\mu\nu}\Box\overline{\phi}\frac{\delta\phi}{\overline{\phi}}\Bigg{]}.$ (15)
 In addition, the scalar field equation at first order (11) turns out:
@@
 Hence, for the weak field limit and by considering the background spacetime metric as the Minkowski metric, one has:
-$\overline{R}=0\ ,\quad T_{\mu\nu}=0\ ,\quad\overline{\phi}=\phi_{0}\ .$ (17)
+$\overline{R}=0\ ,\quad\overline{R}_{\alpha\beta}=0,\quad T_{\mu\nu}=0\ ,\quad\overline{\phi}=\phi_{0}\ .$ (17)
 Then, by definition, $V^{\prime}(\phi_{0})=0$, and in order to satisfy the background solution $V(\phi_{0})=0$ must hold. Hence, the equation for the perturbations (15) reads:
-$\frac{1}{2}\left(\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}h+\Box h\eta_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\eta_{\mu\nu}\right)=\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}-\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ .$ (18)
+$\frac{1}{2}\big{(}\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}$
+$-\nabla_{\nu}\nabla_{\mu}h+\Box h\eta_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\eta_{\mu\nu}\big{)}$
+$=\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}-\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ .$ (18)
 This equation is free from gauge choices and covariant derivatives are actually partial derivatives, as correspond for a flat background metric.
-In order to simplify the equations (18), we can choose the Lorentz Gauge $\partial^{\alpha}h_{\alpha\beta}=0$ to obtain a wave equation.
+In order to simplify the equations (18), we can choose the Lorentz Gauge $\partial^{\alpha}\theta_{\alpha\beta}=0$ to obtain the usual wave equation.
 Note also that the r.h.s.
@@
 By defining the new tensor $\theta_{\mu\nu}=h_{\mu\nu}-\frac{1}{2}h\eta_{\mu\nu}$ the gravitational wave equation yields:
-$\Box\theta_{\mu\nu}=-\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}+\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ .$ (19)
+$\Box\theta_{\mu\nu}=-2\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}+2\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ .$ (19)
 In comparison to GR, now the GWs equations in vacuum are not effectively in vacuum but are sourced by the perturbation of the scalar field.
@@
 By assuming the Lorentz gauge, the equations for the GWs (15) reduce to:
-$\Box\theta_{\mu\nu}=-\frac{2\kappa^{2}}{\phi_{0}}T_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}+\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ ,$ (22)
+$\Box\theta_{\mu\nu}=-\frac{2\kappa^{2}}{\phi_{0}}T_{\mu\nu}-2\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}+2\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ ,$ (22)
 where remind that we have defined $\theta_{\mu\nu}=h_{\mu\nu}-\frac{1}{2}h\eta_{\mu\nu}$. While the equation that relates the scalar field perturbation with the energy-momentum tensor (16) becomes:
@@
 In fact, by combining both equations (22) and (23), the situation becomes clearer:
-$\Box\theta_{\mu\nu}=-\frac{2\kappa^{2}}{\phi_{0}}T_{\mu\nu}+\frac{1}{\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\left(\nabla_{\mu}\nabla_{\nu}T-\eta_{\mu\nu}\Box T\right)\ .$ (24)
+$\Box\theta_{\mu\nu}=-\frac{2\kappa^{2}}{\phi_{0}}T_{\mu\nu}$
+$+\frac{2}{\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\left(\nabla_{\mu}\nabla_{\nu}T-\eta_{\mu\nu}\Box T\right)\ .$ (24)
 Hence, the generation of the GWs does not depend only on the presence of perturbed sources of matter but also on the way the energy-momentum tensor varies, as denoted by the presence of second partial derivatives in the equation (24).
@@
 Nevertheless, in order to keep the analysis as general as possible and study the possible deviations from GR, we are considering that the last two terms might be the same order as the energy-momentum tensor perturbation itself.
-To solve the equation (24), we define the effective energy momentum tensor as:
-$T_{\mu\nu}^{\text{eff}}=\frac{1}{\phi_{0}}T_{\mu\nu}-\frac{1}{2\kappa^{2}\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\left(\nabla_{\mu}\nabla_{\nu}T-\eta_{\mu\nu}\Box T\right)\ ,$ (25)
+To solve the equation (24), we define the effective energy-momentum tensor as:
+$T_{\mu\nu}^{\text{eff}}=\frac{1}{\phi_{0}}T_{\mu\nu}-\frac{1}{\kappa^{2}\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\left(\nabla_{\mu}\nabla_{\nu}T-\eta_{\mu\nu}\Box T\right)\ ,$ (25)
 Then, the GW equation (22) can be rewritten in the standard GR case as:
@@
 $G(x-x^{\prime})=-\frac{1}{4\pi|\vec{x}-\vec{x}^{\prime}|}\delta(x_{ret}^{0}-x^{0\prime})\ .$ (29)
-Here $\vec{x}$ are spatial vectors, $x^{0}=t$ and $x_{ret}^{0}=t_{ret}=t-|\vec{x}-\vec{x}^{\prime}|$ is the retarded time and we are assuming $c=1$.
+Here $\vec{x}$ are spatial vectors, $x^{0}=t$ and $x_{ret}^{0}=t_{ret}=t-|\vec{x}-\vec{x}^{\prime}|$ is the retarded time and we are assuming $c=1$. Then, the solution (27) yields:
+$\theta_{\mu\nu}=4G\int d^{3}x^{\prime}\frac{1}{|x-x^{\prime}|}T_{\mu\nu}^{\text{eff}}(x^{\prime})\ .$ (30)
 By considering an enough large distance $|\vec{x}|$ from the source, we can approximate $|\vec{x}-\vec{x}^{\prime}|=r-\vec{x}^{\prime}\cdot\hat{n}$, where $r=|\vec{x}|$ and $\hat{n}$ is the unit vector pointing to the direction of the GW propagation.
-For simplicity, we work in the transverse-traceless gauge, which by following the standard notation, it is given by:
+Since the volume integral in (30) is taken above the source, the transverse-traceless (TT) gauge can be applied, which by following the standard notation, it is given by:
 $\theta_{ij}^{TT}=\Lambda_{ijkl}\theta^{kl}\ ,$ (31)
@@
 $\Lambda_{ijkl}(\hat{n})=P_{ik}P_{jl}-\frac{1}{2}P_{ij}P_{kl}\ ,\quad P_{ij}=\delta_{ij}-n_{i}n_{j}\ .$ (32)
-Then, by inserting the Green function (29) in the expression (27) and integrating over $x^{0\prime}$, it yields:
+Then, the integral (30) leads to:
 $\theta_{ij}^{TT}=\frac{4G}{r}\Lambda_{ijkl}\int d^{3}x^{\prime}T^{kl}_{\text{eff}}(t-r+\vec{x}^{\prime}\cdot\hat{n},x^{\prime})\ .$ (33)
-The aim now is to perform a multipole expansion and show the dependence of the quadrupole moment for this class of theories.
+Note that inside the source, the TT gauge can not be applied and the following calculations will not be valid what implies important differences, as pointed at the end of the paper.
+Nevertheless, the aim now is to perform a multipole expansion and show the dependence of the quadrupole moment for this class of theories.
 To do so, let’s expand the energy-momentum tensor.
 For enough large distance, we can make the following expansion:
-$T_{kl}^{\text{eff}}(t-\frac{r}{c}+\frac{\vec{x}^{\prime}\cdot\hat{n}}{c},x^{\prime})=T_{kl}^{\text{eff}}(t-r,x^{\prime})+\frac{\partial T_{kl}^{\text{eff}}}{dt}\frac{\vec{x}^{\prime}\cdot\hat{n}}{c}+\frac{\partial^{2}T_{kl}^{\text{eff}}}{dt^{2}}\frac{(\vec{x}^{\prime}\cdot\hat{n})(\vec{x}^{\prime}\cdot\hat{n})}{2c^{2}}+\mathcal{O}\left(\frac{1}{c^{3}}\right)\ ,$ (33)
-where we have recovered the speed of light to show that the expansion is inversely proportional to $c$. Hence, the expression (32) becomes:
-$\theta_{ij}^{TT}=\frac{4G}{r}\Lambda_{ijkl}\left(S^{kl}+\frac{1}{c}n_{m}\dot{S}^{klm}+\frac{1}{2c^{2}}n_{m}n_{p}\ddot{S}^{klmp}+...\right)\ .$ (34)
+$T_{kl}^{\text{eff}}(t-\frac{r}{c}+\frac{\vec{x}^{\prime}\cdot\hat{n}}{c},x^{\prime})=T_{kl}^{\text{eff}}(t-r,x^{\prime})$
+$+\frac{\partial T_{kl}^{\text{eff}}}{\partial t}\frac{(\vec{x}^{\prime}\cdot\hat{n})}{c}+\frac{\partial^{2}T_{kl}^{\text{eff}}}{\partial t^{2}}\frac{(\vec{x}^{\prime}\cdot\hat{n})(\vec{x}^{\prime}\cdot\hat{n})}{2c^{2}}+\mathcal{O}\left(\frac{1}{c^{3}}\right)\ ,$ (34)
+where we have recovered the speed of light to show that the expansion is inversely proportional to powers of $c$. Hence, the expression (33) becomes:
+$\theta_{ij}^{TT}=\frac{4G}{rc^{4}}\Lambda_{ijkl}\bigg{(}S^{kl}+\frac{1}{c}n_{m}\dot{S}^{klm}+$
+$\frac{1}{2c^{2}}n_{m}n_{p}\ddot{S}^{klmp}+...\bigg{)}\ .$ (35)
 Here we have defined the tensors $S^{kl..}$ as follows:
@@
 Then, the leading term in (36) yield:
-$S^{kl}$ $=$ $\int d^{3}x^{\prime}T_{\text{eff}}^{km}\delta_{m}^{l}=\int d^{3}x^{\prime}\frac{\partial}{\partial x^{m\prime}}\left(T_{\text{eff}}^{km}x^{l\prime}\right)-\int d^{3}x^{\prime}\frac{\partial T_{\text{eff}}^{km}}{\partial x^{m\prime}}x^{l\prime}=$ (37)
-$=$ $-\int d^{3}x^{\prime}\frac{\partial T_{\text{eff}}^{km}}{\partial x^{m\prime}}x^{l\prime}=\int d^{3}x^{\prime}\dot{T}_{\text{eff}}^{k0}x^{l\prime}=\int d^{3}x^{\prime}\ddot{T}_{\text{eff}}^{00}x^{k\prime}x^{l\prime}\ .$ (38)
+$S^{kl}$ $=$ $\int d^{3}x^{\prime}T_{\text{eff}}^{km}\delta_{m}^{l}$ (38)
+$=$ $\int d^{3}x^{\prime}\frac{\partial}{\partial x^{m\prime}}\left(T_{\text{eff}}^{km}x^{l\prime}\right)-\int d^{3}x^{\prime}\frac{\partial T_{\text{eff}}^{km}}{\partial x^{m\prime}}x^{l\prime}=$
+$=$ $-\int d^{3}x^{\prime}\frac{\partial T_{\text{eff}}^{km}}{\partial x^{m\prime}}x^{l\prime}=\int d^{3}x^{\prime}\dot{T}_{\text{eff}}^{k0}x^{l\prime}$
+$=$ $\int d^{3}x^{\prime}\ddot{T}_{\text{eff}}^{00}x^{k\prime}x^{l\prime}\ .$
 Here we have used the continuity equation (37) and the volume of integration is taken over the source, such that boundary terms are removed.
@@
 Then, by taking now the expression for the effective energy-momentum tensor (25), the expression for the quadrupole moment tensor is finally obtained:
-$S^{kl}=\frac{1}{\phi_{0}}\frac{\partial^{2}}{\partial t^{2}}\int d^{3}x^{\prime}\left(T^{00}-\frac{1}{\phi_{0}V_{0}^{\prime\prime}}\partial_{j}\partial^{j}T\right)x^{k\prime}x^{l\prime}=\ddot{Q}^{kl}\ ,$ (39)
+$S^{kl}$ $=\frac{1}{\phi_{0}}\frac{\partial^{2}}{\partial t^{2}}\int d^{3}x^{\prime}\left(T^{00}-\frac{1}{\kappa^{2}\phi_{0}V_{0}^{\prime\prime}}\partial_{j}\partial^{j}T\right)x^{k\prime}x^{l\prime}$
+$=\ddot{Q}^{kl}\ ,$ (39)
 where
-$Q^{kl}=\frac{1}{\phi_{0}}\int d^{3}x^{\prime}\left(T^{00}-\frac{1}{\phi_{0}V_{0}^{\prime\prime}}\partial_{j}\partial^{j}T\right)x^{k\prime}x^{l\prime}$ (40)
-is the quadrupole moment for Palatini $f(R)$ gravity. Then, the expression for the gravitational wave (34) at the quadrupole order yields:
-$\left[\theta_{ij}^{TT}\right]_{\text{quad}}=\frac{4G}{rc^{4}}\Lambda_{ijkl}\ddot{Q}^{kl}\ .$ (41)
+$Q^{kl}=\frac{1}{\phi_{0}}\int d^{3}x^{\prime}\left(T^{00}-\frac{1}{\kappa^{2}\phi_{0}V_{0}^{\prime\prime}}\partial_{j}\partial^{j}T\right)x^{k\prime}x^{l\prime}$ (40)
+is the quadrupole moment for Palatini $f(R)$ gravity.
 The first term in the r.h.s.
-of (40) corresponds to the usual one of GR but with an effective gravitational constant given by $G_{eff}=\frac{G}{\phi_{0}}$.
-Nevertheless, the second term turns out a new contribution that just shown up in this class of gravitational theories.
-Such contribution is removed for a homogeneous and isotropic energy-momentum tensor, but might become important in other cases.
-However, even whether the tensor $T^{\mu\nu}$ represents a inhomogeneous and/or anisotropic source, one would expect that variations along different spatial directions might be small enough to become negligible in the weak-field limit.
-Nevertheless, such a term depends on the vacuum expectation value of the scalar field potential and specifically on its second derivative, that can be defined as the scalar field effective mass:
-$m_{eff}^{2}=V^{\prime\prime}(\phi_{0})\ .$ (42)
-Hence, one can conclude that for massive scalar fields, the second contribution in (40) can be neglected while light fields might provide an important contribution to the generation of GWs, even whether the source is nearly homogeneous and isotropic but not completely.
-In addition, at the quadrupole order the energy radiated with the gravitational wave is:
+of (40) corresponds to the usual quadrupole of GR but with an effective gravitational constant given by $G_{eff}=\frac{G}{\phi_{0}}$ while the second one is an extra contribution that arises in this class of gravitational theories.
+However, this second term turns out to vanish in the transverse-traceless gauge, when it is substituted in equation (35).
+Let us show this fact by integrating this term by parts:
+$\int d^{3}x^{\prime}\partial_{j}\partial^{j}Tx^{k\prime}x^{l\prime}$ $=$ $\int d^{3}x^{\prime}\partial_{j}\left(\partial^{j}Tx^{k\prime}x^{l\prime}\right)$
+$-\int d^{3}x^{\prime}\partial^{j}\left(T\partial_{j}\left(x^{k\prime}x^{l\prime}\right)\right)$ $+$ $\int d^{3}x^{\prime}T\partial_{j}\partial^{j}\left(x^{k\prime}x^{l\prime}\right)\ ,$ (41)
+$=$ $2\int d^{3}x^{\prime}T\delta^{kl}\ ,$
+where again the boundary terms are zero because the volume of integration is taken over the source.
+In this way, the second term in the r.h.s.
+of (40) is proportional to $\delta^{kl}$ and vanishes when computing (35), since $\Lambda_{ijkl}\delta^{kl}=0$.
+Therefore, although the GWs propagation equation in Palatini $f(R)$ differs from the one predicted in GR, they are equivalent in the transverse-traceless gauge, at least to quadrupolar order in the multipolar expansion.
+At this point, one can continue the analysis by computing higher order terms in the multipolar expansion.
+In fact, by following a similar procedure as in (38), an expression for $\dot{S}^{klm}$ in terms of $\dddot{T}_{\text{eff}}^{00}$ and $\ddot{T}_{\text{eff}}^{k0}$ can be obtained.
+With this purpose, we use the definition in (36) for the octupolar term and the fact that the energy-momentum tensor is conserved (37).
+However, in this case due to the cyclic behaviour of the integration by parts, we can obtain the following two implicit equations:
+$\dot{S}^{klm}+\dot{S}^{kml}$ $=$ $-\int d^{3}x^{\prime}\ddot{T}_{\text{eff}}^{k0}x^{l\prime}x^{m\prime}\ ,$ (42)
+$\dot{S}^{klm}+\dot{S}^{lmk}+\dot{S}^{mkl}$ $=$ $\frac{1}{2}\int d^{3}x^{\prime}\dddot{T}_{\text{eff}}^{00}x^{k\prime}x^{l\prime}x^{m\prime}\ .$ (43)
+By combining both equations, (42) and (43), an explicit expression for $\dot{S}^{klm}$ is obtained:
+$\dot{S}^{klm}=\frac{1}{2}\int d^{3}x^{\prime}\dddot{T}_{\text{eff}}^{00}x^{k\prime}x^{l\prime}x^{m\prime}+\int d^{3}x^{\prime}\ddot{T}_{\text{eff}}^{m0}x^{l\prime}x^{k\prime}\ .$ (44)
+Here, for convenience, let us split the effective energy-momentum tensor into two terms:
+$T_{\mu\nu}^{\text{eff}}=T_{\mu\nu}^{\text{GR}}+T_{\mu\nu}^{\text{PA}}\ ,$ (45)
+where,
+$T_{\mu\nu}^{\text{GR}}$ $=$ $\frac{1}{\phi_{0}}T_{\mu\nu}\ ,$ (46)
+$T_{\mu\nu}^{\text{PA}}$ $=$ $-\frac{1}{\kappa^{2}\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\left(\nabla_{\mu}\nabla_{\nu}T-\eta_{\mu\nu}\Box T\right)\ .$ (47)
+Note that equation (46) is the usual energy-momentum tensor of GR except for the effective gravitational constant while equation (47) is the contribution of Palatini $f(R)$ theories.
+With that in mind, we can also split (44) into its GR-like part and Palatini $f(R)$ contribution.
+The latter reads:
+$\dot{S}_{\text{PA}}^{klm}=\frac{1}{2}\int d^{3}x^{\prime}\dddot{T}_{\text{PA}}^{00}x^{k\prime}x^{l\prime}x^{m\prime}+\int d^{3}x^{\prime}\ddot{T}_{\text{PA}}^{m0}x^{l\prime}x^{k\prime}\ .$ (48)
+Then, by inserting equation (47) into (48), it leads to:
+$\dot{S}_{\text{PA}}^{klm}$ $=$ $-\frac{1}{\kappa^{2}\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\frac{\partial^{3}}{\partial t^{3}}\bigg{[}\frac{1}{2}\int d^{3}x^{\prime}\partial_{j}\partial^{j}Tx^{k\prime}x^{l\prime}x^{m\prime}$ (49)
+$+$ $\int d^{3}x^{\prime}\partial^{m}Tx^{l\prime}x^{k\prime}\bigg{]}\ .$
+Hence, analogously to (41) and keeping in mind that boundary terms vanish, the expression for $\dot{S}^{klm}$ is reduced to:
+$\dot{S}_{\text{PA}}^{klm}=-\frac{1}{\kappa^{2}\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\frac{\partial^{3}}{\partial t^{3}}\int d^{3}x^{\prime}T\delta^{kl}x^{\prime m}\ .$ (50)
+Whereas the corresponding contraction with the normal vector $n$ yields:
+$n_{m}\dot{S}_{\text{PA}}^{klm}=-\frac{1}{3\kappa^{2}\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\frac{\partial^{3}}{\partial t^{3}}\int d^{3}x^{\prime}T\delta^{kl}n_{m}x^{\prime m}\ .$ (51)
+Recall that the lambda tensor vanishes when contracted with the Kronecker delta, i.e.
+$\Lambda_{ijkl}\delta^{kl}=0$, this implies that the Palatini $f(R)$ contribution to the octupolar term to equation (35) vanishes as well.
+Hence, the octupole order turns out the same as in GR:
+$\dot{S}^{klm}=\frac{1}{2}\int d^{3}x^{\prime}\dddot{T}^{00}x^{k\prime}x^{l\prime}x^{m\prime}+\int d^{3}x^{\prime}\ddot{T}^{m0}x^{l\prime}x^{k\prime}\ .$ (52)
+One might go further in the multipole expansion, but same results will be obtained.
+Hence, in the weak-field limit, the multipolar expansion in Palatini $f(R)$ gravity coincide exactly with the GR predictions.
+In this sense, if one just keeps the lowest order of the expansion (quadrupole order), the expression for the gravitational wave (35) reads:
+$\left[\theta_{ij}^{TT}\right]_{\text{quad}}=\frac{4G}{rc^{4}}\Lambda_{ijkl}\ddot{Q}^{kl}\ ,$ (53)
+where the quadrupole moment (39) reduces to the one of GR:
+$Q^{kl}$ $=\frac{1}{\phi_{0}}\int d^{3}x^{\prime}T^{00}x^{k\prime}x^{l\prime}\ .$ (54)
+Since the energy radiated by the GWs emission is given by:
 $\left[\frac{dE}{dt}\right]_{\text{quad}}\propto\langle\ddot{Q}^{kl}\ddot{Q}^{kl}\rangle\ ,$ (55)
@@
 Then, the full set of field equations are linearised at first order in perturbations of the spacetime metric and the scalar field over an arbitrary background.
-To analyse the weak-field regime, we focus then on studying first order tensor perturbations of the metric in vacuum and also with the presence of weak enough sources.
-For the former case, the usual wave equation of GR is recovered, such that GWs propagation within the Palatini formalism obey the same rules as in GR with respect to polariations and speed of propagation.
-Nevertheless, when assuming the presence of sources and analysing the emission of GWs through retarded Green functions, one finds that the multipoles expansion not only depend on time variations of the energy density, as provided by the 00-component of the energy-momentum tensor but also in spatial derivatives of its trace.
-Hence, for homogeneous and isotropic sources, the same results as in GR are recovered and the lowest order in the multipole expansion (the quadrupole moment) leads to the dependence as in GR.
-However, the presence of inhomogeneous sources introduces inevitable corrections to every order at the expansion and particularly, it modifies the quadrupole moment, leading to a stronger or weaker GWS emission, depending on the source.
-We suggest that for an enough massive scalar field, such corrections are neglected, while light scalar fields are ruled out, since will introduce strong deviations, making the energy radiated by the source too weak or too strong to satisfy observational constraints.
-Next steps suggest that GWs must be analysed as well in strong-field regimes within Palatini-like theories, since the analysis of this paper suggest that while in vacuum, usual perturbation equations of GR are recovered, this changes completely with the presence of sources.
+To analyse the weak-field regime, we focus on studying first order tensor perturbations of the metric in vacuum and also with the presence of weak enough sources.
+For the former case, the usual wave equation of GR is recovered, such that GWs propagation within the Palatini formalism obey the same rules as in GR with respect to polarisations and speed of propagation.
+When assuming the presence of sources and analysing the emission of GWs through retarded Green functions, despite that one finds that the GWs equation is sourced by an effective energy-momentum tensor that depends on derivatives of the trace of the energy-momentum tensor, the multipole expansion only depends on time variations of the energy density, as provided by the 00-component of the energy-momentum tensor.
+Hence, the same results as in GR are recovered and the lowest orders in the multipole expansion (the quadrupole and octupole moments) lead to the same dependence as in GR.
+One expects that higher orders in the multipole expansion will lead behave in the same way.
+A remarkable point is that these results are independent of the form of the gravitational action, such that any $f(R)$ action within the Palatini formalism will lead to the same multipole expansion as in GR.
+Nevertheless, here we have considered just weak sources and assumed a flat background on which the perturbations arise.
+One would expect that for strong-field regimes, the emission of GWs will differ from GR predictions and will depend somehow on the gravitational action $f(R)$ or in other words, on the potential for the scalar field.
+The initial expressions for the quadrupole (39) and octupole moments (49) show an inverse dependence with respect to the effective scalar field mass $m_{eff}^{2}=V^{\prime\prime}(\phi_{0})$ that suggest that in strong field regimes, light scalar field might introduce large deviations with respect to GR predictions whereas for an enough massive scalar field, such corrections might be neglected.
+Within this framework, next steps suggest that GWs must be analysed as well in strong-field regimes for Palatini-like theories, since the analysis of this paper suggest that while in linear theory, usual perturbation equations of GR are recovered, this might change completely with the presence of strong enough sources.
 In addition, the extension of these analysis to arbitrary RBG theories will provide a grateful insight on the understanding of theories within the Palatini formalism.